OpenStudy (anonymous):

find the difference quotient for the following: a. f(x)=x-15 b. f(x)=x^2-x+3

OpenStudy (zzr0ck3r):

((x+h)-15-x+15)/h = (x+h-15-x+15)/h = 1

OpenStudy (zzr0ck3r):

do you see what I did? ((x+h)-f(x))/h b) (x+h)^2-(x+h)+3-(x^2-x+3))/h (x^2+2xh+h^2-x-h+3-x^2+x-3)/h = (2xh+h^2-h)/h = 2x+h-1 then run the limit (not sure if you are asking for this part) 2x-1

OpenStudy (anonymous):

no limit thats all

OpenStudy (zzr0ck3r):

do you see what I am doing to solve these, the biggest problem that students do is forget to distribute the negative through on the - f(x) part.